OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

razor99 (razor99):

what did you mean

OpenStudy (flying_dutchman):

4(x-6) = -2(x-3) | First you have to multiply 4x - 24 = -2x + 6 | Then you have to group the similar terms 6x = 30 | Solve for x x = 30/6 | Simplify x = 5

OpenStudy (anonymous):

thanksss